

********************************************************************
UARS WEEKLY STATUS REPORT - 20 OCT 2000
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible.
We've seen less than 2 arc-minute pointing errors on the SSPP this week
which is getting closer to the typical pointing errors we saw prior to the
star tracker failure. These errors continue to be within our calibrations.
Additionally, we continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
However, level2 processing is halted due to deficiencies in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem and plan to have
a patch by the end of October.
***********
SUSIM
***********
SUSIM continues to gather a normal set of daily solar scans scheduled
for predicted periods of UARS telemetry.
We are preparing to begin the reprocessing the SUSIM UV spectral
irradiances to data version V21 in the next few weeks.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith spectrometers are activate in daylight mode.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in two status meetings this week, one on
Tuesday, October 17 and one on Wednesday, October 18.
The October 17 meeting was held to report the current status
developments concerned with UARS attitude control using the Fine
Sun Sensor (FSS) and the three-axis magnetometer (TAM). The work
is on schedule, but on-board processing of TAM data will not be
available for the October 23 yaw maneuver.
The meeting on October 18 was a working group meeting to review
the script for the October 23 yaw maneuver. This yaw will be
scheduled to occur after the spacecraft passes through 0 beta,
and will be performed using the Fine Sun Sensor (FSS) along with
input from flight dynamics based on TAM data and information from
past yaw maneuvers. The FSS "window" will be enlarged when the
yaw maneuver is completed to accelerate re-acquisition of the
sun.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 07 October 2000 (Orbit 49591, GMT Day 281) through 13
October 2000 (Orbit 49696 , GMT Day 287).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 281/0216 49592 On
HALOE 282/0217 49607 Off
HALOE 282/0256 49608 On
HALOE 283/1739 49632 Off
HALOE 283/1822 49632 On
The beta angle ranged between the angles of 51.0 degrees to 37.7 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
2 degrees on 23 October 2000 (DOY 297).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 to 1.0 degrees C. Additional battery performance data for each
battery on 07 October 2000 and 13 October 2000 are provided in Appendix D.
The clock error ranged from -11.1 to 13.7 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 04
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 36
¥ HALOE = 04
¥ WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/221 - 2287.496937
· Transponder B: 00/244 - 2287.499127
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/272 show the observatory data loss to be 298 hours, 35
minutes, 58.733 seconds (an increase of 5 minutes 58.733 seconds since the last
report period). This is a 0.39 percent data loss which equals a 99.61 percent
data capture for the mission.
The increase of 5 minutes 58.733 seconds occurred over three (3) days:
· 3 minutes 39.944 seconds on Day 281 due to data hits,
· 0 minutes 46.584 seconds on Day 282 due to a data hit,
· 0 minute 58.616 seconds on Day 285 due to data hits.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Testing continues daily to get the TAM and FSS attitude control solution
operational. The UARS testing and training simulator (UTTS) is being used for
all testing. In order to insure the validity of the test the FOT has been
debugging some problems with the simulator attitude models. So far gyro
alignments and propagation errors have been found as well as star tracker
alignment errors and a flight software bug.
MEETINGS
The FOT attended a meeting to discuss UARS attitude control on 10 October 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew has been rescheduled for October 23, 2000 (DOY 297).
This maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing). The maneuver was delayed so that maximum support from Flight
Dynamics, CDHF and the FOT would be present should an anomaly occur.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
_______________________________________________________________________________
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 13 Oct 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
49592 281/0216 00-547 WINDII Power ON
49594 281/0535 00-546 NBTR NOOPS D-281
49594 281/0535 00-548 OBC Clk Adj 0.00 to 20.6 msec/day
49602 281/1736 00-546 NBTR NOOPS D-281
49602 281/1819 00-549 NBTR NOOPS D-282
49607 282/0217 00-550 HALOE Power OFF
49608 282/0256 00-551 HALOE Power ON
49609 282/0537 00-549 NBTR NOOPS D-282
49617 282/1737 00-549 NBTR NOOPS D-282
49618 282/1918 00-552 NBTR NOOPS D-283
49624 283/0441 00-553 OBC Clk Adj 20.6 to 0.00 msec/day
49624 283/0538 00-552 NBTR NOOPS D-283
49631 283/1642 00-554 OBC Attitude Contigency
49632 283/1739 00-552 NBTR NOOPS D-283
49632 283/1739 00-555 HALOE Power OFF
49632 283/1822 00-556 NBTR NOOPS D-284
49632 283/1822 00-557 HALOE Power ON
49639 284/0443 00-556 NBTR NOOPS D-284
49645 284/1421 00-559 OBC Attitude Contigency
49647 284/1740 00-556 NBTR NOOPS D-284
49647 284/1823 00-558 NBTR NOOPS D-285
49653 285/0401 00-558 NBTR NOOPS D-285
49660 285/1422 00-562 OBC Attitude Contigency
49662 285/1742 00-562 NBTR NOOPS D-285
49662 285/1824 00-561 NBTR NOOPS D-286
49659 285/1242 00-560 OBC Clk Adj 0.00 to 20.6 msec/day
49669 286/0543 00-561 NBTR NOOPS D-286
49675 286/1424 00-563 OBC Attitude Contigency
49677 286/1744 00-561 NBTR NOOPS D-286
49677 286/1826 00-564 NBTR NOOPS D-287
49682 287/0124 00-565 OBC Clk Adj 20.6 to 0.00 msec/day
49684 287/0545 00-564 NBTR NOOPS D-287
49691 287/1648 00-566 OBC Attitude Contigency
49692 287/1745 00-564 NBTR NOOPS D-287
49692 287/1745 00-567 OBC Attitude Contigency
49692 287/1827 00-568 NBTR NOOPS D-288
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 13 Oct 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE QB CURRENT MON 49597
HASMIRRT Y-HI SECONDARY MIRROR TEMP 49607
HATBAT Y-HI BOLOMETER AREA TEMP 49607
HASMIRRT Y-HI SECONDARY MIRROR TEMP 49625-6
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 13 Oct 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
49594 281/0546 (07 October) from 0.00 to 20.6 msec/day
49624 283/0444 (09 October) from 20.6 to 0.00 msec/day
49659 285/1256 (11 October) from 0.00 to 20.6 msec/day
49682 287/0127 (13 October) from 20.6 to 0.00 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 13 Oct 2000
06 Oct 2000 - GMT Day 280 Beta = 51.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.62 +0.0/+0.0 14.2 0.000 0.0
2 +16.8/-56.0 +4.83/+3.23 +24.8/-15.2 25.4 1.040 20.5
3 +11.2/-72.8 +3.87/+2.29 +25.2/-15.2 24.8 1.040 20.0
13 Sept 2000 - GMT Day 287 Beta = 37.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.2 0.000 0.0
2 +16.8/-67.2 +5.47/+4.51 +28.8/-15.2 25.0 1.030 21.0
3 +5.6/-140.0 +4.51/+3.23 +28.4/-15.2 25.0 1.030 20.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 13 Oct 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-043 FOT 49643 287/1930 23027 Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-042 FOT see list Oct 00 14170 Generic Late Acq.
00-040 FOT 48854 231/2111 22831 Data Dropout
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/41